i) Describe a Foucault pendulum.

ii) Choose your latitude λ on Earth and write the equation of motion (EOM) for the pendulum.

iii) Use relevant octave script to plot the oscillations both in static and rotating coordinate systems. iv)Discuss when and how, in the actual experiment with pendulum in your lab, you would see the phenomena you see in part iii)


1
Expert's answer
2021-01-27T09:31:04-0500

i) Foucault pendulum is a device which describe about the rotation of earth which is named on the name of French physicist Léon Foucault . In this two iron collars are attached to the cable near the top.




ii)

"mr\theta¨+2r^.\theta˙=2mr˙\Omega \sin\lambda"



iii)